one drain pipe can empty a swimming in 6 hr. Another pipe takes 3 h. if both are used at the same time to drain the pool, what part of the job is completed in 2h? in x h?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

one drain pipe can empty a swimming in 6 hr. Another pipe takes 3 h. if both are used at the same time to drain the pool, what part of the job is completed in 2h? in x h?
One drain pipe can empty in 6 hours
1/6 th. is emptied in 1 hour
Another pipe empties in 3 hours
1/3 rd. is emptied in 1 hour
They both together will empty 1/3+ 1/6 in 1 hour
= 1/2 the pool in 1 hour.
The pool is emptied in 2 hours.
x/2 part of the job is done in x hours



Answer by richwmiller(17219)

h/6+h/3=1
h+2h=6
3h=6
h=2
2/6+2/3=1
in 2 hours the job is 100% completed
in 1 hr 1/2 of the job is completed
if x <2 then x/2 of the job is completed in x hours.